Suppose the market model is described by the following demand and supply functions: Qd= 200-2P+ 4i and Qs= -200 + 3P where i =50 Answer the following questions: a) Calculate the equilibrium price (P) and quantity (Q) and indicate your answers on the graph. [03] b) If the market demand decreases by 50% what will be the new clearing price and quantity for this market? [03] c) Indicate the effect of the changes in (b) on a graph and explain what will happen to market price and quantity.


1
Expert's answer
2022-02-16T10:37:56-0500

At equilibrium P and Q,

"Qd=Qs"

"200-2p+4i=-200+3p"

"400+(4\u00d750)=5p"

"5p=600"

"P=120"

"Q=200-2\u00d7129+4\u00d750"

"Q=160"

B)



C) if "Qd=169×\frac{50}{100}=80"

"Qd=200-2p+4i"

"80=200-2p+200"

"2p=480"

"P=240"

C) The market price will decrease.

The quantity supplied will also decrease.